OpenStudy (anonymous):

Find the exact value of the following. Do NOT use labels in your answer and do NOT round your answer. Find -xy, for x = -75.9 and y = 2.6 -xy = can any body help on this one

OpenStudy (kropot72):

\[\large -xy=-(-75.9\times2.6)\]

OpenStudy (anonymous):

so its -197.34 ??

OpenStudy (kropot72):

Not really. You have correctly multiplied the terms inside the brackets. However your result needs to be multiplied by negative 1.

OpenStudy (kropot72):

\[\large -xy=-(-75.9\times2.6)=-1\times-75.9\times2.6\]

OpenStudy (anonymous):

i just done that and got the same answer

OpenStudy (kropot72):

What is the result of this calculation: \[\large -197.34\times-1=\]

OpenStudy (anonymous):

im using a calculator and i up in that right there and the answer was the same

OpenStudy (kropot72):

I am also using a calculator. However your calculator needs to have a key that can make a number negative in sign. A negative number multiplied by a negative number gives a positive number.

OpenStudy (kropot72):

For example: \[\large -2\times-5=10\]

OpenStudy (kropot72):

@decoda107king Do you follow? The answer to the question that you posted needs to be positive. Remove the negative sign.

OpenStudy (anonymous):

yea i got it lol when u gave the example i cought on but thank u alot

OpenStudy (anonymous):

@kropot72